OpenStudy (anonymous):

\frac{ 3 }{ 4 }k - 5 = \frac{ 1 }{ 4 }k - 1

OpenStudy (anonymous):

\[\frac{ 3 }{ 4 }k - 5 = \frac{ 1 }{ 4 }k - 1\]

OpenStudy (anonymous):